Bond Valuation with Semiannual Payments

Renfro Rentals has issued bonds that have a 6% coupon rate, payable semiannually. The bonds mature in 16 years, have a face value of $1,000, and a yield to maturity of 11%. What is the price of the bonds? Round your answer to the nearest cent.
Answer #1

Face value=1000
Coupon rate=6%, semiannual coupon rate=6%/2=3%
Semiannual coupon payment=(Semiannual coupon rate)*(Face value)=3%*1000=30
Time period=16 years
As the interest is compounded semiannually, the number of periods=16*2=32
Yield to maturity=11%, semiannual yield to maturity=11%/2=0.055

V 1 Face value 1000 2 Payment 30 Number of periods 3 32 Yield to maturity 0.055 ($627.40) 5 Present value= Formula used: PV(V
Hence, the present value of the bond is $627.40